(1) 3x+y=1
-6x-3y=-2
(2) 3x-5y=15
x-2y=8
(3)-2x+y=5
8x-4y=12
(4)2(x+1)=-3y+a
3y-10=-4y
(5) 3(x-2)+2y=5
6y=2x
(6)5x-2y=4
y=-3x+a
(7)4x=5y=0
8(x-1)=10y

Answer by Alan3354(69443)

(1) 3x+y=1
-6x-3y=-2
------
9x + 3y = 3 ---- 1st eqn times 3
-6x -3y =-2
------------------------ Add
3x = 1
x = 1/3
---
3x+y=1
3*(1/3) + y = 1
y = 0
============================
Do the others the same way.
email via the TY note to check your work.
============================
(2) 3x-5y=15
x-2y=8
(3)-2x+y=5
8x-4y=12
(4)2(x+1)=-3y+a
3y-10=-4y
(5) 3(x-2)+2y=5
6y=2x
(6)5x-2y=4
y=-3x+a
(7)4x=5y=0
8(x-1)=10y